OpenStudy (anonymous):

How could you eliminate the fractions to make solving the equation five-twelfths m plus two-twelfths equals eleven-twelfths easier? Divide both sides of the equation by five-twelfths. Divide both sides of the equation by twelve. Multiply both sides of the equation by twelve. Subtract two-twelfths from both sides of the equation. can some please help

OpenStudy (anonymous):

@LukeBlueFive can you help me with this one please

OpenStudy (lukebluefive):

What's the "m" in the equation?

OpenStudy (anonymous):

variable

OpenStudy (lukebluefive):

Oh, okay. First, write out the equation: \[\frac{5}{12}*m + \frac{2}{12} = \frac{11}{12}\]

OpenStudy (anonymous):

okay now what

OpenStudy (lukebluefive):

Next, go through the list of answers and try each one, looking for the one that eliminates the fractions.

OpenStudy (lukebluefive):

You could also ask yourself what operation you could perform to both sides of the equation that would eliminate the fractions, specifically the twelves on the bottom.

OpenStudy (anonymous):

i think it is 3

OpenStudy (lukebluefive):

That's correct!

OpenStudy (anonymous):

thank you for the help again @LukeBlueFive

OpenStudy (lukebluefive):

No problem, anytime.